# Non-unitary enhanced transfer efficiency in quantum walk search on complex networks

### v2

@@ Abstract
The task of finding an element in an unstructured database is known as spatial search and can be expressed as a quantum walk evolution on a graph. In this article, we modify the usual search problem by adding an extra trapping vertex to the graph, which is only connected to the target element. We study the transfer efficiency of the walker to a trapping site, using the search problem as a case study. Thus, our model offers no computational advantage for the search problem, but focuses on information transport in an open environment with a search Hamiltonian. The walker evolution is a mix between classical and quantum walk search dynamics. The balance between unitary and non-unitary dynamics is tuned with a parameter, and we numerically show that depending on the graph topology and the connectivity of the target element, this hybrid approach can outperform a purely classical or quantum evolution for reaching the trapping site. We show that this behavior is only observed in the presence of an extra trapping site, and that depending on the topology, the increase of non-unitary operations can be compensated by increasing the strength of the quantum walk exploration. This compensation comes at the cost of reducing or even suppressing the searching feature of the evolution. In that case, the unitary dynamics will only induce a regular quantum walk behavior. We also relate the optimal hybrid regime to the entropy’s decay rate. As the introduction of non-unitary operations may be considered as noise, we interpret this phenomena as a noisy-assisted quantum evolution.
@@ I Introduction
Quantum walks are the quantum analog of random walks. They are a coherent model of transportation on graphs and a universal model of quantum computation [1, 2], formulated both in continuous and discrete-time. Continuous-time quantum walks (CTQWs) evolve on the space spanned by the vertices of a graph whose structure is encoded in an Hamiltonian [3]. As for discrete-time quantum walks (DTQWs), they require the use of a coin to guide the displacements, which enlarges the Hilbert space of the system [4]. Both formalisms are fundamental quantum computing tools as they serve for quantum simulation of fundamental physics [5, 6, 7, 8, 9] and quantum field theory [10, 11, 12], quantum information processing and quantum algorithms [13, 14]. Among their algorithmic applications, few examples are related to optimization problems [15, 16, 17, 18, 19, 20], quantum state preparation [21, 22, 23], machine learning tasks [24, 25, 26, 27] or graph related problems [28, 29, 30]. Moreover, it has been proved that some DTQWs converge to the Dirac [31, 32, 33, 34, 35] and the Schrödinger [36] equations in their continuous limit.
@@
In this article, we use continuous-time dynamics to tackle a modified version of the search problem for single marked element. We introduce the Stochastic Quantum Walk Search (SQWS) monitored by a weighted Lindbladian. The unitary evolution is induced by Childs and Goldstone’s CTQW search Hamiltonian [42], and non-unitary dissipation is designed to implement a CTRW search dynamics. In addition, we use a trapping sink vertex as an extra dissipative tool, which we only connect to the target vertex of the search with an irreversible transition. Although this is a search problem, we do not quantify performance by the time taken to reach the marked element as a function of graph size as often done, but by the transfer efficiency to reach the trapping site. This metric measures both the amount of information present in the sink and the time needed to reach it. It has already been used as a measure of transport quality in previous works [53, 55, 57, 58], and verified experimentally [54]. In this work, we are interested in a transport problem, not a computational one, and we use the search problem as a case study. The walker starts from a uniform superposition over the vertices of the graph and has to reach the trapping site by moving through a search-driven dynamics guiding it to the target vertex. We numerically show that a mix of unitary and non-unitary operations can outperform a fully coherent or incoherent dynamics for reaching the sink. The balance between unitary and non-unitary operations is controlled with a tunable mixing parameter. The performance of the hybrid regime depends on the graph topology and the connectivity of the target vertex. We show that an increase of non-unitary operations may be compensated by increasing the strength of the quantum walk exploration, at the cost of reducing the importance of the searching oracle that marks the target vertex. Moreover, we show that the addition of non-unitary operations leads to an improvement of performance only in the presence of a sink. Lastly, we relate the best mixing parameter of unitary and non-unitary dynamics to the system’s entropy decay rate. We interpret the hybrid regime as a noisy-assisted quantum evolution as it contains non-unitary operations, even if does not model realistic hardware noise.
@@ II.1.1 Random Walks
The CTRW is a Markov process whose rate matrix is the Laplacian $L=D-A$ of the graph, with $D$ a matrix whose diagonal entries are the degree of each vertex. The state of the walker is described by a probability distribution $\vec{p}$ which is a map from $V$ to probabilities. The time evolution of the walker is:
@@ II.1.2 Quantum Walks
Continuous-time quantum walks are a coherent model of transportation over complex networks [59]. The walker is represented with a quantum state |$\ket{\psi}$ that evolves in an Hilbert space $\mathcal{H}$ spanned by the vertices of the graph. Therefore, the set of vertices $V$ form an orthonormal basis of $\mathcal{H}$. The time evolution of the walker is given by the Schrödinger equation and the Hamiltonian encodes the structure of the graph:
$i\hbar\frac{d}{dt}\ket{\psi(t)}=H\ket{\psi(t)}.$ (3)
@@ II.1.3 Quantum Stochastic Walks
The transition from the classical to the quantum regime can be studied with the QSW framework that enables to interpolate between coherent and incoherent dynamics [45]. The state of the walker is described by a density matrix $\rho=\ket{\psi}\bra{\psi}$ | as the evolution is composed of both unitary and non-unitary operations. The QSWs evolution is driven by a weighted Lindblad master equation with $\omega\in[0,1]$:
@@
where $[\alpha,\beta]=\alpha\beta-\beta\alpha$ and $\{\alpha,\beta\}=\alpha\beta+\beta\alpha$ are the commutator and anti-commutator of operators $\alpha$ and $\beta$. The unitary and non-unitary dynamics are respectively encapsulated by $\mathcal{U}_{H}$ and $\mathcal{D_{L}}$, and the parameter $\omega$ enables to interpolate between them. A wise choice of Lindblad jump operators $\mathcal{L}_{ij}$ can describe a CTRW evolution [45]. Therefore, for well-defined Lindblad jump operators, the CTQW is recovered for $\omega=0$ and the CTRW for $\omega=1$. A linear combination of the two is obtained for other values of $\omega$, leading to a mix between coherent and incoherent dynamics for the walker. It was shown that the introduction of an extra sink vertex in Eq. (4), with $H=A$ and $\mathcal{L}_{ij}=A_{ij}/D_{ij}\ket{i}\bra{j}$ |, leads to an optimal transfer from a set of vertices to this sink vertex when $\omega=0.1$ for several graphs [53].
@@ II.2.1 Random Walk Search
Spatial search expressed as a CTRW evolution consists of a walker spreading over the vertices of a graph until it reaches the marked element $m$. The marked vertex is an absorbing vertex, meaning that the probability of leaving it is zero. Therefore, the Laplacian has to be modified into an absorbing Laplacian $\hat{L}=\frac{D-\hat{A}}{||D-\hat{A}||_{2}}$ where $\hat{A}$ is the adjacency matrix whose $m$-th column is the $m$-th canonical basis vector of $\mathbb{R}^{|V|}$ [60]. The initial state is the uniform probability distribution over the vertices of the graph:
@@ II.2.2 Quantum Walk Search
$H_{m,\gamma}=\gamma L-\ket{m}\bra{m},$ (6)
where $\gamma\in\mathbb{R}^{+}$ determines the strength of the quantum walk exploration induced by $L$ [^1] Note that in the numerical simulations we have normalized the Laplacian so that its eigenvalues are between 0 and 1, thus $L=I-L/\lambda_{\max}$ where $\lambda_{\max}$ is the highest eigenvalue of $L$.. The initial state is the uniform superposition over the vertices of the graph:
|$\ket{\psi(0)}=\frac{1}{\sqrt{|V|}}\sum_{v\in V}\ket{v}.$ (7)
The heart of the CTQWs spatial search algorithm is to find the minimal value of $t$ and the optimal value of $\gamma$ to maximize the success probability $|\bra{m}e^{-itH_{m,\gamma}}\ket{\psi(0)}|^{2}$ of finding the marked element. This algorithm was first shown to offer a quadratic speedup over its classical counterparts for the complete graph, the hypercube and the $d$-dimensional periodic lattice for $d>4$ [42], and to be optimal for a wide family of graphs by Chakraborty et al. [61]. Lastly, it was shown that quantum walk search performances in this framework can be predicted if certain conditions on the spectral properties of the Hamiltonian driving the walk are met [62].
@@ III.1.1 Model
with $\Gamma\in\mathbb{R}^{+}$ the sink rate and $\mathcal{L}_{ij}=-\hat{L}_{ij}\ket{i}\bra{j}$ |. Thus, the coherent and incoherent dynamics respectively produce a quantum and classical random walks search. The initial state is the uniform superposition over the vertices of the graph (excluding the sink vertex) $\rho(0)=\ket{\psi(0)}\bra{\psi(0)}$ | [^2] According to our definitions, $\rho(0)$ evolves in a $|V|$-dimensional Liouville space $\mathcal{B}[\mathcal{H}]$. However, as we add the sink vertex $\phi$ to the graph, its dimension should be $|V|+1$. By abuse of notation, we assume that all the operators defined in Eq. (9) and $\rho(0)$ act on the ($|V|+1$)-dimensional Liouville space spanned by the vertices of the graph and the additional sink vertex $\phi$.. As an illustration, we show in Fig. 1 a graph with the extra sink vertex $\phi$ connected to the target vertex $m$.
@@
$E(\omega,\gamma,t)=\frac{1}{t}\int_{0}^{t}\text{Tr}\left({\mathcal{E}[\rho(\tau)]\ket{\phi}\bra{\phi}}\right)\,\text{d}\tau.$ (10)
@@ III.1.2 Numerical results
We run the SQWS on instances of different graph families and show the results in Fig. 2. We set the time to be linear with the size of the graph as we do not measure performance with computational complexity of reaching the marked element, but with transfer efficiency to the sink. Thus, the value of the maximum evolution time is not as important as for the study of the usual quantum search problem where the optimal performance leads to a quadratic speedup over classical methods. For this reason, we select a large time evolution $t=10|V|$. Moreover, we set the sink rate to $\Gamma=1$. For each graph, we run the SQWS for different fixed values of $\gamma$, which controls the strength of the quantum walk exploration in Eq. (6). For low values of this parameter, the interaction between the Laplacian and the oracle is balanced, but increasing this value implies reducing or even eliminating the action of the oracle, and thus obtaining a unitary dynamic that only induces a simple quantum walk and not a search dynamic towards the marked vertex. Thus, for high values of $\gamma$, the oracle’s action is ignored, eliminating the search dynamic and leading to a free quantum walk. In the following, we use graphs of size $N\in[32,81]$.
@@
We choose fixed values of $\gamma\in\{0.05,0.5,1,2,4,6,$ 8,10,15,20,30,40,50}8,10,15,20,30,40,50\}8 , 10 , 15 , 20 , 30 , 40 , 50 }, taking low and high values to see the impact of the value of this parameter on the transfer efficiency to the sink vertex. We present the main results in Fig. 2 which illustrates the execution of the SQWS on different graphs of 64 vertices, namely, the complete graph $K_{64}$, the 6-dimensional hypercube $Q_{6}$, the cycle graph $C_{64}$, the lollipop graph $L_{32,32}$ and the tadpole graph $T_{32,32}$. The lollipop $L_{M,N}$ and tadpole $T_{M,N}$ graphs are respectively the fusion between a complete or cycle graph of size $M$ with a path of size $N$, we display two instances of these graphs in Fig. 3.
@@
Figure 2: Stochastic Quantum Walk Search (SQWS) performances with different values of $\gamma\in[0.05,50]$ on graphs of 64 vertices, namely, the complete graph ($K_{64}$), the 6-dimensional hypercube graph ($Q_{6}$), the cycle graph ($C_{64}$), the lollipop graph ($L_{32,32}$) and the tadpole graph ($T_{32,32}$). The time evolution is set to $t=640$ units of time. The quantum walk dynamics is recovered for $\omega=0$, the classical random walk for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
We first observe that for all five graphs, when $\gamma=0.05$ the quantum regime, i.e. $\omega=0$, is always the least efficient, with a maximum transfer efficiency of 2%. As we increase the value of the $\gamma$ parameter, we observe different behaviors depending on the observed graph. We point out that for the complete, the hypercube and the cycle graphs, the search behavior will be the same no matter which vertex is marked, because these graphs are vertex-transitive, i.e. their structure does not allow vertices to be distinguished from each other. For the complete graph, we see that $\gamma=0.5$ is sufficient to make the quantum regime efficient, with a transfer efficiency of 80%. The quantum regime reaches its maximum efficiency for $\gamma=2$ with an efficiency of 97%, then increasing $\gamma$ slightly reduces performance to fall back to 90% when $\gamma=50$. The hybrid regime exceeds 70% efficiency as early as $\gamma=0.05$ for $\omega>0.3$, and increasing $\gamma$ does not particularly affect performance. In addition, we note an interesting behavior for the hybrid regime: as the value of $\omega$ increases, so does the transfer efficiency. This is clearly observed with the lowest performance at $\omega=0.1$, corresponding to 90% unitary dynamics and 10% non-unitary. Thus, as we increase the importance of non-unitary dynamics, performance improves. However, the purely quantum regime remains the most efficient. For the hypercube, we observe a behavior quite similar to that of the complete graph, in the sense that the purely quantum regime is the most efficient from a low value of $\gamma$, namely $\gamma=4$, and that increasing $\omega$ induces a performance improvement for the hybrid regime. However, this phenomenon is less significant than for the complete graph, since this time for $\omega=0.1$ we obtain a transfer efficiency of over 60% from $\gamma=2$. However, the hybrid regime for $\omega<0.6]$ is much less efficient than when $\omega\geq 0.6]$ for $\gamma<2$, and then performance stabilizes when $\gamma\geq 2$ for the hybrid regime. The case of the cycle graph is different from the previous ones, as the quantum regime requires a much higher value of $\gamma$ to outperform the hybrid regime. For example, there is a 50% difference in transfer efficiency at $\gamma=4$ between the $\omega=0$ and $\omega=0.1$ regimes. This difference in performance is reduced with increasing $\gamma$, and the $\omega\leq 0.9$ regimes exceed 80% efficiency from $\gamma=50$. The purely quantum regime becomes the most efficient from $\gamma=20$ and reaches a maximum transfer efficiency of 95% at $\gamma=40$. Unlike the complete and hypercube graphs, the classical regime, i.e. $\omega=1$, is not efficient as it only achieves an efficiency of 30%. Furthermore, we observe that for the cycle graph the increase of $\omega$ leads to a reduction in performance unlike in the previous cases. For the lollipop graph, we have run the SQWS on three different vertices that we call complete, shared and path. The complete vertex is an arbitrary vertex of the complete graph and shared is the vertex shared by the complete and path graphs. We also point out that the performance of the SQWS on the path graph is similar to that obtained with the cycle and we present these results in detail in Appendix B. The search results obtained for these two vertices are practically similar, and very different from those for the complete graph seen above. In fact, the introduction of a path graph into the complete graph completely alters performance. Transfer efficiency does not exceed 65% for this graph no matter the marked vertex. By way of comparison, the usual complete graph reaches an efficiency of 97% for $\gamma=2$. The path vertex is located at the extreme end of the path graph, and its search gives different results. Whatever the value of $\gamma$, transfer efficiency to the sink does not exceed 45%. Searching the path vertex is even less efficient than searching the complete and shared vertices, since the 20% efficiency threshold is exceeded at $\gamma=$15 and the 40% threshold at $\gamma=$40. Moreover, the quantum regime does not exceed 12% and the classical one is at 5% efficiency. Lastly, the hybrid regime is more efficient than the quantum one, with a difference of 31% in efficiency when $\gamma=50$ between the quantum case and the hybrid for $\omega\in[0.3,0.6]$. Finally, we ran the SQWS on the tadpole graph on the vertices cycle, shared and path. The vertex cycle is located on the cycle in such a way as to maximize the distance to shared. Unlike the previous case of the lollipop, which was made up of two graphs with very different reactions to the increase in the $\gamma$ parameter, the tadpole graph is a fusion of two graphs with the same reaction to it (the cycle and path graphs). The SQWS behaves approximatively the same way for all three vertices, the hybrid regime performs much better than the quantum regime for certain $\gamma$ values. In particular, we note that, depending on the vertex sought, the value of $\gamma$ required for the quantum regime to be at least as efficient as the hybrid varies. The vertex shared is the one requiring the lowest value of $\gamma$ for this, followed by cycle and finally path. In addition, we observe that for the path vertex, the hybrid regime has more difficulty compensating for the increase in $\omega$ by increasing $\gamma$. Still for this vertex, the quantum regime is not the most efficient even for $\gamma=50$, unlike the other two vertices. Finally, we see that for this graph, the classical regime performs best when searching for the shared vertex.
In this study, we recover the results of Caruso et al. [53, 54] as the value $\omega=0.1$ is critical for graphs where the hybrid regime outperforms the quantum one. In these cases, we observe that the mixing $\omega=0.1$ is the first hybrid regime to outperform quantum dynamics. Moreover, we generalize their result by adding a new parameter $\gamma$ and showing that the topology of the graph plays a major role in the performance of the SQWS. The parameter $\gamma$ plays a fundamental role as it allows to control the strength of quantum walk exploration. However, we see that an increase of $\gamma$ reveals a new spectrum of the hybrid behavior as the increase of $\omega$ can be compensated with an increase of $\gamma$. However, an increase in $\gamma$ tends to reduce the importance of the oracle that marks the searched vertex and thus produce dynamics more centered on a quantum walk free rather than a search quantum walk. Generally speaking, the SQWS behaves in a number of interesting ways, depending on the graph topology and the target vertex connectivity. The first behavior is the possibility for a hybrid regime to outperform quantum and classical dynamics for this modified search problem. We also note that in some cases, a low-noise hybrid regime (low value of $\omega$) is less efficient than a noisier hybrid regime. This phenomenon is clearly observed for the complete graph, but also in a lighter way for the hypercube and lollipop (for the complete and shared vertices) graphs. Finally, we see that increasing the value of $\gamma$ can compensate for the increase in the interpolation parameter $\omega$. In other words, increasing the importance of graph exploration for the walker, and thus neglecting the oracle that enables the search, offsets the increase in the importance of incoherent operations in the system.
@@
Intuitively, we can use metrics in an attempt to understand why the SQWS behaves differently on different graphs, and even on different vertices belonging to the same graph. A useful global graph metric is its density, with the densest complete graph serving as a reference with a density of 1. Then, two interesting local metrics for the target vertex are its eccentricity, i.e. the longest of the shortest paths to reach that vertex from any vertex in the graph, and its degree centrality, which indicates how connected the vertex is in the graph. A centrality of 1 means that the vertex is connected to all the others, and 0 to none. Interestingly, density and eccentricity are equal quantities for vertex transitive graphs, i.e. graphs whose structure does not allow vertices to be distinguished from each other. We show all these characteristics for all the graphs on which we have run the SQWS on Table 1. We can see that all the graphs where the hybrid out regime performs the quantum regime for certain values of $\gamma$ have a high eccentricity, i.e. which scales in $\mathcal{O}(N)$. Among these graphs (cycle, path, maze, lollipop, tadpole) we also observe that the higher the eccentricity and the lower the centrality of the target vertex, then the purely quantum regime requires a higher value of $\gamma$ to outperform the hybrid. Furthermore, we see that the lower the eccentricity and centrality, the lower the value of $\gamma$ required for the quantum regime to perform. Finally, it seems that the higher the density and centrality, the less an increase in $\gamma$ can compensate for the introduction of incoherent dynamics into the search.
@@
Lollipop $L_{M,N}$ $M$+$N\!=\!32$+$32$ 0.2619 complete 0.4920 $33=N+1=\mathcal{O}(N)$

### v3

@@ Abstract
The task of finding an element in an unstructured database is known as spatial search and can be expressed as a quantum walk evolution on a graph. In this article, we modify the usual search problem by adding an extra trapping vertex to the graph, which is only connected to the target element. We study the transfer efficiency of the walker to a trapping site, using the search problem as a case study. Thus, our model offers no computational advantage for the search problem, but focuses on information transport in an open environment with a search Hamiltonian. The walker evolution is a mix between classical and quantum walk search dynamics. The balance between unitary and non-unitary dynamics is tuned with a parameter, and we numerically show that depending on the graph topology and the connectivity of the target element, this hybrid approach can outperform a purely classical or quantum evolution for reaching the trapping site. We show that this behavior is only observed in the presence of an extra trapping site, and that depending on the topology and a tunable parameter controlling the strength of the oracle, a hybrid regime composed of 90% coherent dynamics can lead to either the highest or worst transfer efficiency to the trapping site. We also relate the performance of an hybrid regime to the entropy’s decay rate. As the introduction of non-unitary operations may be considered as noise, we interpret this phenomena as a noisy-assisted quantum evolution.
@@ I Introduction
Quantum walks are the quantum analog of random walks. They are a coherent model of transportation on graphs and a universal model of quantum computation [1, 2], formulated both in continuous and discrete-time. Continuous-time quantum walks (CTQWs) evolve on the space spanned by the vertices of a graph whose structure is encoded in an Hamiltonian [3]. As for discrete-time quantum walks (DTQWs), they require the use of a coin to guide the displacements, which enlarges the Hilbert space of the system [4]. Both formalisms are fundamental quantum computing tools as they serve for quantum simulation of fundamental physics [5, 6, 7, 8, 9] and quantum field theory [10, 11, 12], quantum information processing and quantum algorithms [13, 14]. Among their algorithmic applications, few examples are related to optimization problems [15, 16, 17, 18, 19, 20, 21], quantum state preparation [22, 23, 24], machine learning tasks [25, 26, 27, 28, 29] or graph related problems [30, 31, 32]. Moreover, it has been proved that some DTQWs converge to the Dirac [33, 34, 35, 36, 37] and the Schrödinger [38] equations in their continuous limit.
@@
In this article, we use continuous-time dynamics to tackle a modified version of the search problem for single marked element. We introduce the Stochastic Quantum Walk Search (SQWS) monitored by a weighted Lindbladian. The unitary evolution is induced by Childs and Goldstone’s CTQW search Hamiltonian [44] where we parametrize the oracle instead of the walk generator, and non-unitary dissipation is designed to implement a CTRW search dynamics. In addition, we use a trapping sink vertex as an extra dissipative tool, which we only connect to the target vertex of the search with an irreversible transition. Although this is a search problem, we do not quantify performance by the time taken to reach the marked element as a function of graph size as often done, but by the transfer efficiency to reach the trapping site. This metric measures both the amount of information present in the sink and the time needed to reach it. It has already been used as a measure of transport quality in previous works [55, 57, 59, 60], and verified experimentally [56]. In this work, we are interested in a transport problem, not a computational one, and we use the search problem as a case study. The walker starts from a uniform superposition over the vertices of the graph and has to reach the trapping site by moving through a search-driven dynamics guiding it to the target vertex. We numerically show that a mix of unitary and non-unitary operations can outperform a fully coherent or incoherent dynamics for reaching the sink. The balance between unitary and non-unitary operations is controlled with a tunable mixing parameter. The performance of the hybrid regime depends on the graph topology and the connectivity of the target vertex. We show that a low-noise regime can outperform others regimes only for graphs with low density, high eccentricity and low degree centrality for the target vertex. Moreover, we show that the addition of non-unitary operations leads to an improvement of performance only in the presence of a sink. Lastly, we relate the performance of a mixing regime to the system’s entropy decay rate. We interpret the hybrid regime as a noisy-assisted quantum evolution as it contains non-unitary operations, even if does not model realistic hardware noise.
@@ II.1.1 Random Walks
The CTRW is a Markov process whose rate matrix is the Laplacian L=DAL=D-A of the graph, with $D$ a matrix whose diagonal entries are the degree of each vertex. The state of the walker is described by a probability distribution $\vec{p}$ which is a map from $V$ to probabilities. The time evolution of the walker is:
@@ II.1.2 Quantum Walks
[…]
@@ II.1.3 Quantum Stochastic Walks
The transition from the classical to the quantum regime can be studied with the QSW framework that enables to interpolate between coherent and incoherent dynamics [47]. The state of the walker is described by a density matrix $|\rho=\ket{\psi}\bra{\psi}$ as the evolution is composed of both unitary and non-unitary operations. The QSWs evolution is driven by a weighted Lindblad master equation with $\omega\in[0,1]$:
@@
where $\alpha,\beta]=\alpha\beta-\beta\alpha$ and $\{\alpha,\beta\}=\alpha\beta+\beta\alpha$ are the commutator and anti-commutator of operators $\alpha$ and $\beta$. The unitary and non-unitary dynamics are respectively encapsulated by $\mathcal{U}_{H}$ and $\mathcal{D_{L}}$, and the parameter $\omega$ enables to interpolate between them. A wise choice of Lindblad jump operators $\mathcal{L}_{ij}$ can describe a CTRW evolution [47]. Therefore, for well-defined Lindblad jump operators, the CTQW is recovered for $\omega=0$ and the CTRW for $\omega=1$. A linear combination of the two is obtained for other values of $\omega$, leading to a mix between coherent and incoherent dynamics for the walker. It was shown that the introduction of an extra sink vertex in Eq. (4), with $H=A$ and $|\mathcal{L}_{ij}=A_{ij}/D_{ij}\ket{i}\bra{j}$, leads to an optimal transfer from a set of vertices to this sink vertex when $\omega=0.1$ for several graphs [55].
@@ II.2.1 Random Walk Search
[…]
@@ II.2.2 Quantum Walk Search
$|,H_{m,\gamma}=L-\gamma\ket{m}\bra{m}$, (6)
where $\gamma\in\mathbb{R}^{+}$ determines the strength of the oracle and $L$ is the graph Laplacian [^1] Note that in the numerical simulations we have normalized the Laplacian so that its eigenvalues are between 0 and 1, thus $L=I-L/\lambda_{\max}$ where $\lambda_{\max}$ is the highest eigenvalue of $L$.. Note that in the literature the search Hamiltonian is usually defined as $H_{\text{search}}=\gamma^{-1}H_{m,\gamma}$ and can be recovered from $H_{m,\gamma}$ with proper time rescaling $t=t_{\text{search}}\gamma^{-1}$. However, we parametrize the oracle instead of the Laplacian to control the strength of the oracle instead of the hopping rate of the walk. The initial state is the uniform superposition over the vertices of the graph:
[…]
@@ III.1.1 Model
with $\Gamma\in\mathbb{R}^{+}$ the sink rate and $|\mathcal{L}_{ij}=-\hat{L}_{ij}\ket{i}\bra{j}$. Thus, the coherent and incoherent dynamics respectively produce a quantum and classical random walks search. The initial state is the uniform superposition over the vertices of the graph (excluding the sink vertex) $|\rho(0)=\ket{\psi(0)}\bra{\psi(0)}$ [^2] According to our definitions, $\rho(0)$ evolves in a $|V|$-dimensional Liouville space $\mathcal{B}[\mathcal{H}]$. However, as we add the sink vertex $\phi$ to the graph, its dimension should be $|V|+1$. By abuse of notation, we assume that all the operators defined in Eq. (9) and $\rho(0)$ act on the ($|V|+1$)-dimensional Liouville space spanned by the vertices of the graph and the additional sink vertex $\phi$.. As an illustration, we show in Fig. 1 a graph with the extra sink vertex $\phi$ connected to the target vertex $m$.
@@
[…]
@@ III.1.2 Numerical results
We run the SQWS on instances of different graph families and show the results in Fig. 2. We set the time to be linear with the size of the graph as we do not measure performance with computational complexity of reaching the marked element, but with transfer efficiency to the sink. Thus, the value of the maximum evolution time is not as important as for the study of the usual quantum search problem where the optimal performance leads to a quadratic speedup over classical methods. For this reason, we select a large time evolution $t=10|V|$. Moreover, we set the sink rate to $\Gamma=1$. For each graph, we run the SQWS for different fixed values of $\gamma$, which controls the strength of the oracle in Eq. (6). For $\gamma=0$, the coherent evolution only induces a free quantum walk exploration, and setting $\gamma>0$ impacts the strength of the oracle marking the target vertex. In the following, we use graphs of size $N\in[32,81]$.
@@
We choose fixed values of $\gamma\in[0,30]$, taking low and high values to see the impact of the value of this parameter on the transfer efficiency to the sink vertex. We present the main results in Fig. 2 which illustrates the execution of the SQWS on different graphs, namely, the complete graph $K_{64}$, the 6-dimensional hypercube $Q_{6}$, the cycle graph $C_{64}$, the path graph $P_{65}$, a maze graph [^4] Maze generation can be easily done using Depth-First Search (DFS) on a grid. Once the maze is created, each cell is considered as a vertex, and two vertices are adjacent if there is no wall between their respective cells. ($M_{73}$) and the tadpole graph $T_{32,32}$. The tadpole $T_{M,N}$ graph is the fusion between a cycle graph of size $M$ with a path of size $N$, we display an instance of this graph [^5] We also display an instance of the lollipop $L_{M,N}$ graph which is a fusion between a complete graph of size $M$ with a path of size $N$, and we run the SQWS on it in Appendix. B). in Fig. 3.
@@
Figure 2: Stochastic Quantum Walk Search (SQWS) performances with different values of $\gamma\in[0,30]$ on graphs of 64 vertices, namely, the complete graph ($K_{64}$), the 6-dimensional hypercube graph ($Q_{6}$), the cycle graph ($C_{64}$), the path graph ($P_{65}$), a maze graph ($M_{73}$) and the tadpole graph ($T_{32,32}$). The time evolution is set to $t=640$ units of time. The quantum walk dynamics is recovered for $\omega=0$, the classical random walk for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.
We first point out that for the complete, the hypercube and the cycle graphs, the search behavior will be the same no matter which vertex is marked, because these graphs are vertex-transitive, i.e. their structure does not allow vertices to be distinguished from each other.
For the complete graph, we observe that all regimes $\omega$ achieve at least 50% transfer efficiency to the sink when $\gamma\leq 2$. We see that for the purely quantum case, i.e. $\omega=0$, increasing the oracle strength with $\gamma$ improves transfer efficiency, reaching a maximum of 97% for $\gamma=1$. Then, when $\gamma>2$ is reached, transfer efficiency begins to decrease, reaching 0.02% for $\gamma=30$. Interestingly, we see that even when $\gamma=0$ which corresponds to the free quantum walk with no oracle, the walker already achieves a transfer efficiency of 87%. We also observe that the classical regime, i.e. $\omega=1$, achieves the highest transfer efficiency with 89% for $\gamma=0$, which corresponds to a free quantum walk in the coherent regime, and does not induce a searching dynamics. Moreover, we note that variations in the parameter $\gamma$ do not seem to affect the hybrid regime when $\omega>0.1$, and furthermore, among the noisy regimes, we see that a higher value of $\omega$ results in better transfer to the sink. Thus, among the hybrid regimes, the worst transfer is obtained for $\omega=0.1$ and the quantum regime produces the worst transfer from $\gamma\geq 5$. For the hypercube, we obtain results similar to those for the complete graph, except that the maximum transfer efficiency achieved is lower, approximately 80% for $\omega=0$. Furthermore, the performance of the purely quantum regime reaches its maximum of 84% around $\gamma\approx 0.4$ and decreases sharply from $\gamma\approx 1$ to reach 0.01 for $\gamma=30$. For the hybrid regime, we observe the same phenomenon as for the complete graph, i.e. transfer efficiency increases with $\omega$, and variations in $\gamma$ only affect the hybrid regime from $\omega\geq 0.3$ onwards. For the other four graphs presented, we obtain different results. For the path graph we have run the SQWS on two different vertices that we call border and center, which are respectively placed at the extremity and center of the graph. For the maze, the target vertex corresponds to the exit and finally for the tadpole graph we chose three different vertices, cycle, shared and path, their locations are shown on Fig. 3 on a smaller instance. First, the quantum regime is not the most efficient for any value of $\gamma$. Furthermore, the maximum transfer efficiency is low compared to the complete graph and the hypercube, as it does not exceed 50%. The interesting phenomenon for these graphs is that, unlike in the previous cases, the regime around $\omega=0.1$ is the most efficient for all values of $\gamma$ tested. There is a clear and significant performance gain in this region when the dynamics consist of 90% coherent evolution and 10% non-unitary operations. For these graphs, the transfer to the sink is greatly improved by adding slight noise to the unitary dynamics.
In this study, we recover part of the results of Caruso et al. [55, 56] as their is an interesting phenomena around $\omega=0.1$ for specific graphs where the low-noise hybrid regime outperforms all the others. However, we only observe this phenomena for sparse graphs, and we obtain the opposite effect for the others, i.e. a significant drop in performance towards the $\omega=0.1$ regime compared to the other regimes. Therefore, the topology of the graph plays a major role in the performance of the SQWS and the parameter $\gamma$ mostly affect the quantum regime, i.e. $\omega=0$. In the quantum case, we observe a sharp decrease in performance when a critical value of $\gamma$ is reached. When this parameter continues to increase, we obtain a transfer of almost zero for the purely quantum regime, because the force of the oracle is too predominant over the exploration of the graph. Generally speaking, the SQWS behaves in a number of interesting ways, depending on the graph topology and the target vertex connectivity. The first behavior is the possibility for a low-noise hybrid regime to outperform quantum, classical and other hybrid dynamics for this modified search problem. We only observe this phenomenon for the cycle, path, maze and tadpole graphs. The second is the opposite behavior, where a low-noise hybrid regime is the worst for transfer efficiency in the context of low values of $\gamma$. This phenomenon is clearly observed for the complete graph and for the hypercube. Moreover, we see that for graphs where the quantum regime leads to a good transfer efficiency to the sink, when a critical value of $\gamma$ is reached, performance declines very rapidly, and thus a hybrid regime is more effective.
@@
Intuitively, we can use metrics in an attempt to understand why the SQWS behaves differently on different graphs, and even on different vertices belonging to the same graph. A useful global graph metric is its density, with the densest complete graph serving as a reference with a density of 1. Then, two interesting local metrics for the target vertex are its eccentricity, i.e. the longest of the shortest paths to reach that vertex from any vertex in the graph, and its degree centrality, which indicates how connected the vertex is in the graph. A centrality of 1 means that the vertex is connected to all the others, and 0 to none. Interestingly, density and eccentricity are equal quantities for vertex transitive graphs, i.e. graphs whose structure does not allow vertices to be distinguished from each other. We show all these characteristics for all the graphs on which we have run the SQWS on Table 1. The common feature of cycle, path, maze and tadpole graphs is that they all have very low density, and their vertices have high eccentricity and low degree centrality. These three characteristics appear to be necessary to observe the performance gain around $\omega=0.1$ that outperforms all other regimes. Furthermore, we note that for center vertex of star and wheel graphs (see Appendix B), performance for $\omega=0.1$ does not decrease significantly as it does for other graphs, which may indicate that low density favours low noise regime performance around $\omega=0.1$. We also see that for the center vertex of the wheel graph, performance is slightly lower for low values of $\gamma$ compared to the star graph, indicating that lower density and degree centrality favour low-noise regimes. Thus, the phenomenon observed by Caruso et al. [55, 56] seems to depend on the topology of the graph and is observed in this study for graphs with low density, poorly connected neighborhood, and which require long paths to reach a vertex from the target vertex.
@@
Lollipop $L_{M,N}$ MM+N=32N\!=\!32+3232 0.2619 complete 0.4920 $33=N+1=\mathcal{O}(N)$